# Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex Optimization

### v2

@@ Abstract
In this paper, we concentrate on decentralized optimization problems with nonconvex and nonsmooth objective functions, especially on the decentralized training of nonsmooth neural networks. We introduce a unified framework to analyze the global convergence of decentralized stochastic subgradient-based methods. We prove the global convergence of our proposed framework under mild conditions, by establishing that the generated sequence asymptotically approximates the trajectories of its associated differential inclusion. Furthermore, we establish that our proposed framework covers a wide range of existing efficient decentralized subgradient-based methods, including decentralized stochastic subgradient descent (DSGD), DSGD with gradient-tracking technique (DSGD-T), and DSGD with momentum (DSGD-M). In addition, we introduce the sign map to regularize the update directions in DSGD-M, and show it is enclosed in our proposed framework. Consequently, our convergence results establish, for the first time, global convergence of these methods when applied to nonsmooth nonconvex objectives. Preliminary numerical experiments demonstrate that our proposed framework yields highly efficient decentralized subgradient-based methods with convergence guarantees in the training of nonsmooth neural networks.
Keywords: Nonsmooth optimization, decentralized stochastic subgradient method, conservative field, differential inclusion, Lyapunov function.
@@ 1 Introduction
Here for any $i=1,\ldots,d$, let $f_{i}({\bm{x}}):=\frac{1}{|\mathcal{S}_{i}|}\sum_{l\in\mathcal{S}_{i}}F_{i}({\bm{x}};{\bm{s}}_{l})$, where $\mathcal{S}_{i}$ denotes the local dataset of agent $i$ and $\{{\bm{s}}_{l}:l\in\mathcal{S}_{i}\}$ represents the samples within dataset $\mathcal{S}_{i}$. Each $F_{i}(\cdot,{\bm{s}}_{l}):\mathbb{R}^{n}\rightarrow\mathbb{R},l\in\mathcal{S}_{i}$ is nonconvex and possibly nonsmooth, and exclusively known to the agent $i$. ${\bm{x}}_{i}$ refers to the local variables of the agent $i$. For convenience, we define the sum function $f$ as
$f({\bm{x}}):=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}})$ (1.1)
Moreover, $\mathtt{E}$ denotes the set of edges of the graph $\mathtt{G}$ and represents communication links between agents, in the sense that all the agents communicate solely with their immediate neighbors within the graph $\mathtt{G}$.
@@ 1.1 Existing works on decentralized stochastic optimization
The decentralized optimization problem (DOP) has wide applications in data science and machine learning [1, 2], leading to increasing demands for efficient optimization methods. Motivated by stochastic gradient descent (SGD), the decentralized stochastic gradient descent (DSGD) [3, 4] is one of the most fundamental methods for solving (DOP). When we assume the differentiability of $\{f_{i}\}$, the update scheme of DSGD can be expressed as
@@
where the subset $\mathcal{B}_{i,k}\subseteq\mathcal{S}_{i}$ refers to the mini-batch selection of the samples. Therefore, for any $i\in[d]:=\{1,\ldots,d\}$, the term $\frac{1}{|\mathcal{B}_{i,k}|}\sum_{l\in\mathcal{B}_{i,k}}\nabla F_{i}({\bm{x}}_{i,k};{\bm{s}}_{l})$ refers to the stochastic gradient of $f_{i}$ at ${\bm{x}}_{i,k}$, In addition, we define $\xi_{i,k+1}:=\frac{1}{|\mathcal{B}_{i,k}|}\sum_{l\in\mathcal{B}_{i,k}}\nabla F_{i}({\bm{x}}_{i,k};{\bm{s}}_{l})-\nabla f_{i}({\bm{x}}_{i,k})$ can be viewed as the corresponding evaluation noise. Moreover, ${\bm{W}}\in\mathbb{R}^{d\times d}$ represents the mixing matrix corresponding to the graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$. That is, ${\bm{W}}$ is symmetric, doubly stochastic, and ${\bm{W}}(i,j)=0$ whenever $(i,j)\notin\mathtt{E}$ (see Section 2.2 for details).
For better understanding of the existing approaches, we give a unified notation system. We denote the nested local variables ${\bm{X}}_{k}$, evaluation noises $\Xi_{k+1}$, and noiseless gradients at $k$-th iteration ${\bm{G}}_{k}$ as
@@
${\bm{G}}_{k}:=$ $[\nabla f_{i}({\bm{x}}_{1,k}),\ldots,\nabla f_{d}({\bm{x}}_{d,k})],$
Then for any $i\in[d]$, we denote the mapping $\Phi^{\mathrm{GD}}_{i}:=\{\nabla f_{i}\}$ to characterize the local updates in the $i$-th agent in the DSGD method. With this notation, the update scheme of DSGD in (1.2) can be reformulated as follows:
${\bm{X}}_{k+1}={\bm{X}}_{k}{\bm{W}}-\eta_{k}({\bm{G}}_{k}+\Xi_{k+1})$ (1.4)
@@
Based on DSGD, a wide range of efficient decentralized methods have been developed for solving (DOP) under the assumption of differentiability of its objective functions. For example, [5, 6] propose the DSGD with momentum (DSGD-M) by integrating the heavy-ball momentum [7] into the DSGD. Moreover, some recent works introduce auxiliary variables within DSGD to track the stochastic gradient of the objective function in each agent, such as the DSGD with gradient-tracking technique (DSGD-T) [8, 9], etc.
To provide theoretical guarantees for these decentralized methods with continuously differentiable objective functions, [4] provides the first theoretical analysis for the convergence rate of DSGD. Moreover, [10, 11, 12, 13, 14] further explore the convergence properties of the DSGD, especially regarding network topology and data heterogeneity. Additionally, [5, 15, 16] analyze the global convergence of the different versions of DSGD-M they proposed, while [17, 18, 9, 19] show the convergence properties of DSGD-T for continuously differentiable objectives.
Furthermore, some existing works develop decentralized proximal gradient methods for (DOP), under the assumption that $f_{i}$ is composed of a differentiable term $g_{i}$ and a convex regularization term $r_{i}$ for all $i\in[d]$. In this context, the objective function is weakly convex and thus Clarke regular. For example, [10] introduces Prox-DGD based on the update scheme of DGD, [20] introduces NEXT based on DGD-T and successive convex approximation techniques, while [21] proposes PG-EXTRA derived from EXTRA. The authors all establish the global convergence of their respective algorithms. Moreover, [22, 23, 24, 25] extend the convergence analysis of Prox-DGD and NEXT from the deterministic case to the stochastic case under different assumptions. Additionally, [26] improves the convergence analysis for DSGD-T and proves its convergence for weakly convex objectives.
Although achieving nice theoretical results, all the aforementioned works are capable of handling, at most, non-smooth but Clarke regular [27] objective functions in (DOP). However, as discussed in [28], the nonsmooth activation functions, including ReLU and leaky ReLU, are popular building blocks for modeling neural networks. More importantly, in the training of neural networks, the widely employed ReLU activation function leads to the nonsmoothness of the loss function. As a result, in the decentralized training of various popular neural networks, the objective functions $\{f_{i}:1\leq i\leq d\}$ in (DOP) are usually nonsmooth without Clarke regularity. All these works can not easily extend their convergence theories to the decentralized training of nonsmooth neural networks. How to establish convergence properties for decentralized stochastic subgradient-based methods under nonconvex and nonsmooth settings remains unexplored.
1.2 Stochastic subgradient-based methods for nonsmooth nonconvex optimization
In the training of nonsmooth neural networks, one major challenge comes from computing the (stochastic) subgradients of the loss function. In a wide range of deep learning packages, the subgradients of the loss function are computed by the automatic differentiation (AD) algorithms. Utilizing the chain rule, these algorithms recursively construct generalized subgradients by composing the Jacobians of each block. However, since the chain rule fails for Clarke subdifferential [27], the outputs of the AD algorithms may not be contained in the Clarke subdifferential of the loss function. To characterize the outputs of AD algorithms in differentiating nonsmooth neural networks, [29, 28] introduce the class of path-differentiable functions, which is general enough to cover a wide range of objective functions in real-world applications, including semi-algebraic functions, semi-analytic functions, and functions whose graphs are definable in some o-minimal structures [29]. For path-differentiable functions, [28] introduces the concept of the conservative field as a generalization of their Clarke subdifferential, and shows that the chain rule holds for conservative fields. As a result, although the results of AD algorithms may not be constrained in the Clarke subdifferential of the loss function of a nonsmooth neural network, these results are contained in a specific conservative field of that loss function. Therefore, we employ the concept of the conservative field as a generalization of the Clarke subdifferential to characterize the outputs of AD algorithms throughout our paper.
To analyze the convergence properties of stochastic subgradient-based methods, the ordinary differential equation (ODE) approaches are developed [30, 31, 32, 29], which describe the convergence of the iterates in a subgradient method through the asymptotic behavior of the dynamics of corresponding differential inclusion (i.e., a nonsmooth generalization of ODE). Then, based on the concepts of path-differentiable functions and conservative field, a wide range of stochastic subgradient-based methods are developed for nonsmooth nonconvex optimization problems within a single agent (i.e., (DOP) with $d=1$), while enjoying convergence guarantees from ODE approaches. [29, 33, 28] study the global convergence of SGD method. Moreover, [34, 35, 36] introduce the SGD with heavy-ball momentum and establish its global convergence. In addition, [37] proposes the inertial Newton method (INNA) and shows its global convergence in training nonsmooth neural networks. Furthermore, [38] proves the global convergence of the AdaGrad-Norm, and [39] provides convergence guarantees for a family of ADAM methods. In addition, [40, 41] discuss stochastic subgradient-based methods for optimization problems over manifolds.
Significant progress has been made in studying the convergence properties of stochastic subgradient-based methods within a single agent. This motivates us to explore decentralized stochastic subgradient-based methods under nonconvex and nonsmooth settings based on the above-mentioned ODE approaches. Compared with single-agent settings, establishing the convergence properties of decentralized stochastic subgradient methods is non-trivial, as the update schemes of all the agents are coupled through local communication under decentralized settings. Although [42] establishes the global convergence of DSGD for nonsmooth functions utilizing the ODE approach, their analysis heavily relies on the condition $\partial(\sum_{i\in[d]}f_{i})=\sum_{i\in[d]}\partial f_{i}$. This assumption presumes the validity of the chain rule for $\{f_{i}\}$ in (DOP). However, the chain rule does not hold for the Clarke subdifferential of general nonsmooth functions, hence limiting their analysis to Clarke regular functions [27]. In the presence of limited existing results towards the convergence of decentralized subgradient-based methods in minimizing nonsmooth path-differentiable functions, developing a unified framework for establishing convergence properties for these methods is of great importance in decentralized optimization.
@@ 1.3 Contributions
The main contributions of our paper are summarized as follows:
  • A novel framework for decentralized stochastic subgradient-based methods
    Motivated by the update scheme of DSGD (1.4) that characterized by the mappings $\{\Phi^{\mathrm{GD}}_{i}\}$, we first introduce the following unified framework for decentralized stochastic subgradient methods,
    ${\bm{Z}}_{k+1}={\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}).$ (DSM)
    Here, ${\bm{Z}}_{k}=[{\bm{z}}_{1,k},\ldots,{\bm{z}}_{d,k}]\in\mathbb{R}^{m\times d}$ denotes the nested local variables, ${\bm{H}}_{k}\in\mathbb{R}^{m\times d}$ refers to the noiseless nested update directions, $\Xi_{k+1}\in\mathbb{R}^{m\times d}$ refers to the evaluation noises, and $\eta_{k}>0$ denotes the step-size. Moreover, we introduce set-valued mappings $\{\Phi_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m},i\in[d]\}$ to characterize the relationship between ${\bm{Z}}_{k}$ and ${\bm{H}}_{k}$, in the sense that $\frac{1}{d}{\bm{H}}_{k}{\bm{1}_{d}}$ is an approximated evaluation of $\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}}_{i,k})$. We demonstrate that, with particular choices of the set-valued mappings $\{\Phi_{i}\}$, the framework (DSM) is general enough to enclose a wide range of decentralized subgradient-based methods, including DSGD, DSGD-M, and DSGD-T.
    We prove that under mild conditions, $\lim_{k\rightarrow+\infty}$ $\left\|{\bm{z}}_{i,k}-{\bm{z}}_{j,k}\right\|=0$ holds for any $i,j\in[d]$, and the sequence $\left\{\boldsymbol{z}_{i,k}\right\}$ asymptotically approximates the trajectories of the differential inclusion
    $\frac{\mathrm{d}{\bm{z}}}{\mathrm{d}t}\in-\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}})\right).$ (1.5)
    When the differential inclusion (1.5) admits Lyapunov function $\psi$ and stable set $\mathcal{A}$, we prove that any cluster point of $\left\{{\bm{z}}_{i,k}\right\}$ lies in $\mathcal{A}$ under mild conditions based on the existing ODE approaches.
  • Convergence guarantees for decentralized stochastic subgradient-based methods
    Based on our proposed framework, we provide convergence guarantees for DSGD, DSGD-M, and DSGD-T. Furthermore, we also propose a decentralized variant of SignSGD [43] denoted as DSignSGD, which can be enclosed by (DSM). As there is no existing work on the convergence properties of DSGD-M, DSGD-T, and DSignSGD in solving nonsmooth nonconvex optimization problems, our results fill the gap between their practical implementations and theoretical guarantees. Preliminary numerical experiments show the efficiency of analyzed methods and exhibit the superiority of our proposed method DSignSGD in comparison with DSGD and DSGD-M, hence demonstrating the flexibility and great potential of our framework.

### v3

# Convergence of Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex functions

@@ Abstract
In this paper, we focus on the decentralized stochastic subgradient-based methods in minimizing nonsmooth nonconvex functions without Clarke regularity, especially in the decentralized training of nonsmooth neural networks. We propose a general framework that unifies various decentralized subgradient-based methods, such as decentralized stochastic subgradient descent (DSGD), DSGD with gradient-tracking technique (DSGD-T), and DSGD with momentum (DSGD-M). To establish the convergence properties of our proposed framework, we relate the discrete iterates to the trajectories of a continuous-time differential inclusion, which is assumed to have a coercive Lyapunov function with a stable set $\mathcal{A}$. We prove the asymptotic convergence of the iterates to the stable set $\mathcal{A}$ with sufficiently small and diminishing step-sizes. These results provide first convergence guarantees for some well-recognized of decentralized stochastic subgradient-based methods without Clarke regularity of the objective function. Preliminary numerical experiments demonstrate that our proposed framework yields highly efficient decentralized stochastic subgradient-based methods with convergence guarantees in the training of nonsmooth neural networks.
Keywords: Nonsmooth optimization, decentralized optimization, stochastic subgradient-based method, random reshuffling, with-replacement sampling, conservative field, Lyapunov function.
@@ 1 Introduction
Here, for each $i\in[d]:=\{1,2,\ldots,d\}$, ${\bm{x}}_{i}$ refers to the local variable of agent $i$. Moreover, the set of nodes $\mathtt{V}=[d]:=\{1,\ldots,d\}$ represents the set of agents, while the set of edges $\mathtt{E}\subseteq\mathtt{V}\times\mathtt{V}$ represents the set of communication links. An edge $(i,j)\in\mathtt{E}$ refers to that agents $i$ and $j$ are neighbors and can communicate with each other. Furthermore, for any $i\in[d]$, $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is a locally Lipschitz continuous (possibly nonconvex and nonsmooth) cost function exclusively known to the agent $i$ and takes the empirical-risk formulation [1]. More precisely, for any $i\in[d]$, there exists $F_{i}:\mathbb{R}^{n}\times\Omega\to\mathbb{R}$ such that
$f_{i}({\bm{x}}):=\frac{1}{|\mathcal{S}_{i}|}\sum_{s_{l}\in\mathcal{S}_{i}}F_{i}({\bm{x}};{\bm{s}}_{l}),$ (1.1)
where ${\bm{s}}_{l}$ is a random data vector supported on a local finite set $\mathcal{S}_{i}$. In real-world applications, the size of $\mathcal{S}_{i}$ is often sufficiently large, in the sense that computing the exact (sub)gradient of $f_{i}$ is very costly, which makes stochastic (sub)gradient-based methods inevitable. Problem (DOP) has wide applications in machine learning [2, 3], signal processing [4, 5] and control [6], etc.
@@ 1.1 Existing works on decentralized stochastic optimization
Motivated by stochastic gradient descent (SGD), the decentralized stochastic gradient descent (DSGD) [7, 8, 9] is one of the most fundamental methods for solving (DOP). In DSGD, each agent locally averages its model parameters with those of its neighbors and updates along the direction of the negative local stochastic gradient, which can be expressed as
@@
where the subset $\mathcal{B}_{i,k}\subseteq\mathcal{S}_{i}$ refers to the mini-batch selection of the samples. Therefore, for any $i\in[d]$, the term $\frac{1}{|\mathcal{B}_{i,k}|}\sum_{l\in\mathcal{B}_{i,k}}\nabla F_{i}({\bm{x}}_{i,k};{\bm{s}}_{l})$ refers to the stochastic gradient of $f_{i}$ at ${\bm{x}}_{i,k}$, In addition, we define $\xi_{i,k+1}:=\frac{1}{|\mathcal{B}_{i,k}|}\sum_{l\in\mathcal{B}_{i,k}}\nabla F_{i}({\bm{x}}_{i,k};{\bm{s}}_{l})-\nabla f_{i}({\bm{x}}_{i,k})$ as the corresponding evaluation noise. ${\bm{W}}\in\mathbb{R}^{d\times d}$ represents the mixing matrix corresponding to the graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$. That is, ${\bm{W}}$ is symmetric, doubly stochastic, and ${\bm{W}}(i,j)=0$ whenever $(i,j)\notin\mathtt{E}$ (see Section 2.2 for details).
Based on DSGD, a wide range of variants have been developed for solving (DOP). Some works introduce auxiliary variables within DSGD to track the global stochastic gradient of the objective function, such as GNSD [10] and DSGT [11]. To accelerate convergence, some recent works [12, 13] propose the decentralized momentum SGD by integrating the heavy-ball momentum [14] into DSGD. Moreover, [15] combines the stochastic gradient tracking method with momentum acceleration. In addition, [16, 17] introduce random reshuffling (RR) [18] sampling techniques (i.e., randomly samples an index $i_{k}$ of $[N]$ without replacement at each step $k$ in agent $i$, ensuring that each index appears exactly once per epoch.) into DSGD and DSGT, and provide theoretical evidence that RR outperforms with-replacement sampling. A series of studies [9, 19, 20, 21, 22] have established theoretical guarantees for these decentralized methods with continuously differentiable nonconvex objective functions.
Much attention has also been directed towards decentralized nonsmooth optimization. Some existing works, such as ProxDGD [23], NEXT [24], and PG-EXTRA [25] develop decentralized proximal gradient type methods for (DOP), under the assumption that $f_{i}$ is composed of a differentiable term $g_{i}$ and a convex regularization term $r_{i}$ for all $i\in[d]$. In this context, the objective function is weakly convex and thus is Clarke regular. Several existing works [26, 27, 28, 29] extend the convergence analysis of Prox-DGD and NEXT from the deterministic case to the stochastic case, under additional assumptions regarding boundedness and heterogeneity. Additionally, [30] proposes stoDPSM for $\rho$-weakly convex objectives and shows the global convergence while employing the Moreau envelope for measuring the stationarity. [31] improves the convergence analysis for DSGT and proves its convergence for weakly convex objectives. From the view of distributed gradient flow, [32] establishes the global convergence of DSGD for Clarke regular function with decreasing consensus parameter. The detailed comparisons between the existing results on decentralized SGD-based methods in nonsmooth nonconvex optimization are listed in Table 1.
Despite achieving nice theoretical results, all the aforementioned works are only capable of handling nonsmooth but (Clarke) regular [33] objective functions in (DOP), while some of these works evaluate the efficiency of their developed methods in the decentralized training of neural networks. However, nonsmooth activation functions, including ReLU and leaky ReLU, have served as popular building blocks for modeling neural networks, resulting in loss functions that are nonsmooth and lack Clarke regularity. To the best of our knowledge, [34] is the only recent work that investigates the convergence of decentralized SGD-based method without requiring Clarke regularity. They propose a decentralized first-order and zero-order online algorithm called Me-DOL, but it requires several strong assumptions, such as global Lipschitz function and $L^{2}$ bounded stochastic subgradients. As a result, how to establish convergence properties for decentralized stochastic subgradient-based methods in nonsmooth, nonconvex settings remains an area for further exploration.
Method Update scheme Step-sizes Conditions on $f_{i}$ Conditions on noises
stoDPSM [30] GD Square summable, $\eta_{k+1}/\eta_{k}\to 1$ $\rho$-weakly convex, bounded subgradient WRS, bounded second moment
S-NEXT [26] GT Square summable Proximal, weakly convex, bounded subgradient WRS, bounded noise
ProxGT [28] GT Constant Proximal, weakly convex, mean-squared smoothness WRS, bounded variance
DGF [32] GD $\eta_{k}=O(k^{-p}),p\in(0,1]$ Clarke regular, weakly coercive, decreasing consensus parameter WRS, bounded variance
ME-DOL [34] GD $\eta_{k}=\Theta(\frac{1}{\sqrt{K}}),k\in[K]$ Global Lipschitz continuous WRS, bounded variance, bounded second moment
Our work (DSM) GD/GT /GD-M $\eta_{k}=o(1/\log(k))$ Path-differentiable, Coercive WRS, conditionally controlled noises
Our work (DSM) GD/GT /GD-M diminishing $\eta_{k}$ Path-differentiable, Coercive RR
Table 1: A brief comparison of existing decentralized SGD-based methods for nonsmooth nonconvex optimization. Here, “GD”, “GT” and “GD-M” are abbreviations of “gradient descent”, “gradient tracking” and “gradient descent with momentum technique”; “WRS” and “RR” are abbreviations of “with-replacement sampling” and “random reshuffling”; The term “conditionally controlled noise” means that the evaluation noise is bounded conditional on $\mathcal{F}_{k}=\sigma(\{{\bm{x}}_{i,l},l\leq k,i\in[d]\})$. The concept ”path-differentiable” is defined in Definition 2.6.
1.2 Existing works on stochastic subgradient-based methods for nonsmooth nonconvex optimization
In the training of nonsmooth neural networks, one major challenge comes from computing the (stochastic) subgradients of the loss function. In a wide range of deep learning packages, such as TensorFlow, PyTorch, and JAX, automatic differentiation (AD) algorithms have become the default choice for computing subdifferential. Utilizing the chain rule, AD algorithms recursively construct the so-called “subgradients” through the composition of Jacobians of each block of the neural network, yet they neglect the underlying nonsmoothness. As the chain rule fails for loss functions lacking Clarke regularity, the so-called “subgradient” may not necessarily be contained in its Clarke subdifferential (see examples in [35]). This limitation consequently renders the theoretical analyses of existing literature inapplicable to such functions.
To characterize the behavior of AD algorithms, [36] introduces the concept of path-differentiable functions, which constitute a broad function class, including semi-algebraic functions, semi-analytic functions, and functions whose graphs are definable in some o-minimal structures [37]. More importantly, the set of path-differentiable functions is general enough to cover almost all loss functions in neural network tasks, encompassing most of the loss functions without Clarke regularity. The study [36] further demonstrates that when applied to path-differentiable loss functions, the outputs of AD algorithms are contained within a conservative field, which is a generalization of the Clarke subdifferential. Conservative field admits chain rules for path-differentiable functions, thereby preserving several good properties. For a comprehensive discussion of this concept, readers are directed to Section 2.3.
Based on the concepts of path-differentiable functions and conservative field, some recent works [36, 38, 39, 40, 41, 42, 43, 44] leverage the ordinary differential equation (ODE) approach [45, 46, 47, 48] to study the behavior of stochastic subgradient-based methods in the non-distributed setting. Although these studies have achieved significant progress in establishing convergence properties for stochastic subgradient-based methods, the theoretical results of many approaches cannot avoid infinitely many spurious critical points due to the gap between the conservative field and the Clarke subdifferential [49, 50]. Moreover, extending existing results to the multi-agent scenario is non-trivial, as the update schemes of all the agents are coupled through local communication under decentralized settings. Given the convergence properties that need improvement and limited existing results in the multi-agent scenario, developing a unified framework for establishing convergence properties for decentralized stochastic subgradient-based methods is of great importance and worth exploring.
1.3 A general framework for decentralized stochastic subgradient-based methods
We define the collection of the local variables, evaluation noises and noiseless gradients in (1.2) by matrix ${\bm{X}}_{k}$, $\Xi_{k+1}$, and ${\bm{H}}_{k}$, respectively:
@@ 1.3 A general framework for decentralized stochastic subgradient-based methods
${\bm{H}}_{k}:=$ $[\nabla f_{1}({\bm{x}}_{1,k}),\ldots,\nabla f_{d}({\bm{x}}_{d,k})].$
Then for any $i\in[d]$, we denote $\Phi^{\mathrm{GD}}_{i}:{\bm{x}}\mapsto\{\nabla f_{i}({\bm{x}})\}$ as a singleton-valued mapping from $\mathbb{R}^{n}$ to $2^{\mathbb{R}^{n}}$, which characterize the noiseless update directions in the $i$-th agent in the DSGD method. With these notations, the update scheme of DSGD in (1.2) can be reformulated as follows:
${\bm{X}}_{k+1}={\bm{X}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})$ (1.4)
@@
In light of the reformulated update scheme (1.4), we consider the following unified framework for decentralized stochastic subgradient-based methods,
${\bm{Z}}_{k+1}={\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}).$ (DSM)
Here, ${\bm{Z}}_{k}=[{\bm{z}}_{1,k},\ldots,{\bm{z}}_{d,k}]\in\mathbb{R}^{m\times d}$ denotes the collection of local variables (e.g. model parameters, momentum, etc.), ${\bm{H}}_{k}\in\mathbb{R}^{m\times d}$ refers to the collection of noiseless update directions, $\Xi_{k+1}\in\mathbb{R}^{m\times d}$ refers to the collection of evaluation noises, and $\eta_{k}>0$ denotes the step-size.
Moreover, the noiseless update directions in (DSM) is characterized by a family of locally bounded graph-closed set-valued mappings $\{\Phi_{i}:i\in[d]\}$ that maps $\mathbb{R}^{m}$ to the collections of subsets of $\mathbb{R}^{m}$. More precisely, there exists a sequence of nonnegative real numbers $\{\delta_{k}\}$ such that
Here, $\Phi_{i}^{\delta}({\bm{x}}):=\{{\bm{y}}\in\mathbb{R}^{n}:\exists{\bm{z}},\text{ s. t. }\left\|{\bm{z}}-{\bm{x}}\right\|<\delta,\inf_{{\bm{w}}\in\Phi_{i}({\bm{z}})}\left\|{\bm{y}}-{\bm{w}}\right\|<\delta\}$.  [^1] The superscript is used to denote expansion only when applied to set-valued mappings. Therefore, it is easy to verify that (DSM) encloses DSGD method in (1.2).
@@ 1.4 Contributions
The main contributions of our paper can be summarized as follows.
We propose a unified framework (DSM) that encompasses a variety of decentralized stochastic subgradient-based methods. We verify that decentralized stochastic subgradient descent (DSGD), DSGD with gradient-tracking technique (DSGD-T), and DSGD with momentum (DSGD-M) all fit into the framework (DSM). In particular, a decentralized variant of SignSGD [51], denoted as DSignSGD, is also enclosed by (DSM).
We establish the convergence guarantees for our framework (DSM). By introducing the following differential inclusion
$\frac{\mathrm{d}{\bm{z}}}{\mathrm{d}t}\in-\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}})\right),$ (1.6)
that admits a coercive Lyapunov function $\psi$ and stable set $\mathcal{A}$, we prove that, under mild conditions with controlled noises, the sequence $\{{\bm{Z}}_{k}\}$ generated by (DSM) is uniformly bounded. Then we prove that the iterates $\{{\bm{z}}_{i,k}\},i\in[d]$ asymptotically reach consensus and approximate the trajectories of the differential inclusion (1.6), hence converging to the stable set $\mathcal{A}$ under sufficiently small and diminishing step-size.
Based on our theoretical results of the proposed framework (DSM), we prove that DSGD, DSGD-M, DSGD-T and DSignSGD achieve global convergence in the random reshuffling case and high-probability global convergence in the with-replacement sampling case. We also show that with random initialization, DSGD, DSGD-M and DSignSGD can avoid the spurious critical points in the sense of conservative field in minimizing path-differentiable functions. Although these algorithms are widely used in the training of nonsmooth neural networks, existing theoretical results are limited to Clarke regular objective functions. Our theoretical findings first fill the gap in solving Problem (DOP) with objectives lacking Clarke regularity.
Preliminary numerical experiments show the efficiency of analyzed methods and exhibit the superiority of our proposed method DSignSGD in comparison with DSGD and DSGD-M, hence demonstrating the flexibility and great potential of our framework.